OpenStudy (anonymous):

Given that b is a real number, for what values of b is the equation true? A) no values of b B) some values of b C) all values of b D) not enough information to make a decision

OpenStudy (anonymous):

this is the equation \[\frac{ 3b+9 }{ 3}=b+3\]

OpenStudy (yrelhan4):

LHS=(3b+9)/3 =3b/3+9/3 =b+3=RHS So C.

OpenStudy (yrelhan4):

Got it?

OpenStudy (anonymous):

yah thanks

OpenStudy (yrelhan4):

You're welcome.

OpenStudy (anonymous):

why you put lhs

OpenStudy (anonymous):

rhs

OpenStudy (yrelhan4):

Left hand side=lhs Right hand side=rhs I just proved that whats written on lhs is equal to rhs. So that basically means it will be that for all values of b.

OpenStudy (anonymous):

ohhhhh